# On sensitivity of nucleus deformation on final-state flow harmonics

### v1

@@ I Introduction
Also, analyses indicated [36, 37] the importance of the multipole deformations. Besides $\beta_{2}$, which measures the quadrupole, such deformation is also characterized by $\beta_{3}$ and $\beta_{4}$, namely, the hexapole and hexadecapole parameters. These parameters are relevant for the multipole expansion of the Woods-Saxon distribution for non-spherical nuclei. In literature, the values of these parameters typically vary significantly in different studies [38, 39, 40, 41, 42].
@@ II.3 Decoupling
As the system evolves hydrodynamically, it eventually cools down, and hadronization, known as decoupling, takes place. We use the THERMINATOR 2, a Monte Carlo approach, in the CHESS code to implement the hadrons decoupling using the Cooper-Frye prescriptions [64, 65]. The Cooper-Frye formula describes the hadronic decoupling as follows
@@
where $d\Sigma_{\mu}$ is an element of the freezeout hypersurface and $f(x,p)$ gives the distribution function, which can be properly modified to take into account the effect of the bulk and shear viscosities.
@@ II.4 The united code: CHESS
These separated files are then returned to “aux” and serve as part of the input files for $T_{R}ENT_{o}$, vHLLE, and THERMINATOR 2. The IC is generated by $T_{R}ENT_{o}$ using the parameters governed by input file I. Subsequently, vHLLE is invoked for the relativistic hydrodynamical evolution. The results obtained from $T_{R}ENT_{o}$ together with the input file part II will be fed to vHLLE as the parameters.
Then, the output from vHLLE will be combined with input file III, which is used as input to THERMINATOR 2 to perform the hadronization. Finally, the postscript is utilized to evaluate the relevant observables, such as particle spectrum and collective flow. The calculations are carried out on an event-by-event fluctuating basis.
@@ II.5 Model setup and calibration
Last but not least, to ensure that we have sufficient statistical power for the numerical analysis, tentative simulations are carried out for U2delimited-⟨⟩2$\langle$ 2$\rangle$ 2 ⟩+U2delimited-⟨⟩2$\langle$ 2$\rangle$ 2 ⟩ collisions at $\sqrt{s_{NN}}=193$ GeV while varying the number of events (NE). The flow harmonics evaluated for different NE are presented in Fig. 5. In the left panel of Fig. 5, the differential elliptic flow $v_{2}$ and its standard deviation are shown as a function of transverse momentum. Also, the dependence of average elliptic flow on the NE is analyzed in the right panel of Fig. 5.
@@ III Numerical results and discussions
Figure 7: Differential elliptic flow $v_{2}$ using the centrality 0-6% for Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV with different degrees of nucleus deformation. The numerical results obtained by the CHESS code are compared with the STAR data [66]. Top-Left: The results obtained by varying quadrupole deformation parameter $\beta_{2}$, the specific parameters are given in Tab. 1. Top-Right: The results obtained by varying hexadecapole deformation parameter $\beta_{4}$, the specific parameters are given in Tab. 1. Bottom: The results obtained by varying the overall deformations of the Au nucleus. For all the plots, the red triangles are used to represent the results for symmetric Au nuclei. The error bars indicate the standard deviations.
It is observed that the initial geometrical deformations do affect the resulting elliptic flow. As shown in the bottom panel of Fig. 7, as the overall deformation increases, the resulting differential flow decreases, pulling the simulation closer to the data. The relation between the size of the deformation and the resulting impact to the elliptic flow is primarily linear. Nonetheless, the model could not faithfully reproduce the data at large transverse momenta. By tuning individual deformation parameters, the above results mostly hold. Specifically, the elliptic flow generally becomes more suppressed, particularly at the region with larger transverse momenta, as $\beta_{2}$ or $\beta_{4}$ increases. At smaller $p_{T}$, on the other hand, the dependence of elliptic flow on these parameters is complex. An explicit study on the dependence of $v_{2}$ on the deformation parameters will be presented below in Fig. 10, where the average flow will be analyzed.
A similar procedure is also carried out for various types of U+U collisions. In Fig. 8, we present the flow harmonics for the three Uranium nuclei while tuning the respective parameters regarding the quadrupole and hexadecapole deformations. The numerical results obtained by CHESS are also compared to the STAR data [66]. Generally, the initial geometric deformation plays a role in the final elliptic flow. Unlike $\beta_{2}$, the differential elliptic flow mainly increases with increasing $\beta_{4}$, which is apparent particularly for the region with large transverse momentum.
@@
Figure 8: Differential elliptic flow $v_{2}$ for U+U collisions at $\sqrt{s_{NN}}=193$ GeV with centrality 0-6% and different degrees of nucleus deformation. The numerical results obtained by the CHESS code are compared with the STAR data [66] for Au+Au collisions at 200 GeV. Top-Left: The results obtained by varying quadrupole deformation parameter $\beta_{2}$ for U. Top-Right: The results obtained by varying hexadecapole deformation parameter $\beta_{4}$ for U. Middle-Left: The results obtained by varying quadrupole deformation parameter for U2. Middle-Right: The results obtained by varying hexadecapole deformation parameter $\beta_{2}$ for U3. Bottom-left: The results obtained by varying hexadecapole deformation parameter $\beta_{4}$ for U3. Bottom-right: The results obtained for symmetric U+U, U2+U2, and U3+U3 collisions. The specific parameters are given in Tab. 1 for all the plots. The error bars indicate the standard deviations.
@@
By comparing the results of Au+Au and U+U collisions, one seems to draw somewhat different observations about the effect of deformation on the collective flow. On the one hand, for Au+Au collisions, we have observed that the increase of deformation has, by and large, suppressed the elliptic flow $v_{2}$, pushing the simulation results closer to the experimental data. On the other hand, a rather opposite effect has been observed in U+U collisions. Regarding the magnitude of the effect, one finds that the elliptic flow in U+U collisions is more sensitive to the deformation.
@@
Figure 9: Elliptic flow $v_{2}$ comparison for Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV and different U+U collisions at $\sqrt{s_{NN}}=193$ GeV. Top-Left: Symmetric Au nuclei compared to symmetric and deformed U nuclei in terms of $\beta_{2}$. Top-Right: Symmetric Au nuclei compared to symmetric and deformed U2 nuclei. Middle-Left: Symmetric Au nuclei compared to symmetric and deformed U3 nuclei in terms of $\beta_{2}$. For all the plots, the red triangles are used to represent the results for symmetric Au nuclei. Middle-Right: Symmetric Au nuclei compared to symmetric and deformed U nuclei in terms of $\beta_{4}$. Bottom: Symmetric Au nuclei compared to symmetric and deformed U3 nuclei in terms of $\beta_{4}$. The error bars indicate the standard deviations.
@@
For Au+Au collisions, it is observed that the elliptic flow is rather sensitive when the deformation is small. When compared with Fig. 7, it seems that the somewhat outstanding value of $\bar{v_{2}}$ is primarily due to the contribution from the anisotropy of high $p_{T}$ particles. Here, the average is taken among all the emitted particles and on an event-by-event basis. As the deformation increases, the impact on the elliptic flow is not monotonic, which eventually becomes convergent. For the present study, we found that the value $\beta_{2}\approx-0.26$ gives the average elliptic flow that is in reasonable agreement with the experimental data. As a comparison, Giacalone [40] found a relatively smaller value $\beta_{2}\approx-0.15$, by considering only the quadrupole deformation. We attribute such a difference mostly to that of the hydrodynamic model and IC.
@@ IV Concluding remarks
Our numerical results prompt a comparison between Au+Au and U+U nuclear collisions, revealing distinct behaviors. In the case of Au+Au collisions, the simulations for both symmetric and deformed nuclei give reasonable results regarding the experimental data from the STAR Collaboration, as shown in Fig. 7. The results for deformed nuclei stay closer to the data. On the other hand, as shown in Fig. 10, the transition from symmetric to deformed nuclei marks a significant shift in average elliptic flow, $\overline{v_{2}}$, owing to the contributions from high transverse momentum $p_{T}\sim 2$ GeV. For symmetric nuclei, $\overline{v_{2}}$ significantly overestimates the data, and as deformation increases, the average elliptic flow converges to the experimental data. Conversely, for U+U nucleus collisions, numerical simulations diverge further from the STAR data as deformation factors increase. These results are explicitly illustrated in Fig. 8, where the elliptic flow increases in response to the deformation factors. The computation of the average elliptic flow, presented in Fig. 10, supports the linear behavior described in Eq. (7) that is extensively documented in the literature.
In the literature, the studies regarding the effect of initial geometric deformation are mainly motivated to answer the following question: whether such deformation is observable and which deformation factor, $\beta_{2}$ or $\beta_{4}$, predominantly influences collective flow? From the present study, the answer remains elusive, aligned with the results obtained by other authors [50, 38, 39, 40]. Nonetheless, for collisions of almost symmetric Au nuclei, the resulting elliptic flow is found to be somewhat sensitive to the initial quadrupole and hexadecapole deformations. In particular, further investigation is needed to understand why $v_{2}$ in Au+Au collisions is particularly sensitive to the deformation factor $\beta_{2}$, unlike in U+U collisions.

### v2

@@ I Introduction
Also, analyses indicated [36, 37] the importance of the multipole deformations. Besides $\beta_{2}$, which measures the quadrupole, such deformation is also characterized by $\beta_{3}$ and $\beta_{4}$, namely, the octupole and hexadecapole parameters. These parameters are relevant for the multipole expansion of the Woods-Saxon distribution for non-spherical nuclei. In literature, the values of these parameters typically vary significantly in different studies [38, 39, 40, 41, 42].
@@ II.3 Decoupling
As the system evolves hydrodynamically, it eventually cools down, and hadronization, known as decoupling, takes place. We use the THERMINATOR 2 [28], a Monte Carlo approach, in the CHESS code to implement the hadrons decoupling using the Cooper-Frye prescriptions [64, 65]. The Cooper-Frye formula describes the hadronic decoupling as follows
@@
where $d\Sigma_{\mu}$ is an element of the freezeout hypersurface and $f(x,p)$ gives the distribution function, which can be properly modified to take into account the effect of the bulk and shear viscosities. Besides hadronization, the model also takes into account the two- and three-body decays of resonances. Although explicit space-time evolution of hadronic degree of freedom will not be considered under the current theoretical framework, we understand that the main results on anisotropy drawn in this study are largely unaffected.
@@ II.4 The united code: CHESS
These separated files are then returned to “aux” and serve as part of the input files for $T_{R}ENT_{o}$, vHLLE, and THERMINATOR 2. The IC is generated by $T_{R}ENT_{o}$ using the parameters governed by input file I. Subsequently, vHLLE is invoked for the relativistic hydrodynamical evolution. The results obtained from $T_{R}ENT_{o}$ together with the input file part II will be fed to vHLLE as the parameters. Then, the output from vHLLE will be combined with input file III, which is used as input to THERMINATOR 2 to perform the hadronization. Via T Finally, the postscript is utilized to evaluate the relevant observables, such as particle spectrum and collective flow. The calculations are carried out on an event-by-event fluctuating basis.
@@ II.5 Model setup and calibration
Last but not least, to ensure that we have sufficient statistical power for the numerical analysis, tentative simulations are carried out for U2delimited-⟨⟩2$\langle$ 2$\rangle$ 2 ⟩+U2delimited-⟨⟩2$\langle$ 2$\rangle$ 2 ⟩ collisions at $\sqrt{s_{NN}}=193$ GeV while varying the number of events (NE). The flow harmonics evaluated for different NE are presented in Fig. 5. In the left panel of Fig. 5, the differential elliptic flow $v_{2}$ and its standard error are shown as a function of transverse momentum. Also, the dependence of average elliptic flow on the NE is analyzed in the right panel of Fig. 5.
@@ III Numerical results and discussions
Figure 7: Differential elliptic flow $v_{2}$ using the centrality 0-6% for Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV with different degrees of nucleus deformation. The numerical results obtained by the CHESS code are compared with the STAR data [66]. Top-Left: The results obtained by varying quadrupole deformation parameter $\beta_{2}$, the specific parameters are given in Tab. 1. Top-Right: The results obtained by varying hexadecapole deformation parameter $\beta_{4}$, the specific parameters are given in Tab. 1. Bottom: The results obtained by varying the overall deformations of the Au nucleus. For all the plots, the red triangles are used to represent the results for symmetric Au nuclei. The error bars indicate the standard errors.
It is observed that the initial geometrical deformations do affect the resulting elliptic flow. As shown in the bottom panel of Fig. 7, the resulting differential flow slightly increases as the overall deformation increases. The relation between the size of the deformation and the resulting impact on the elliptic flow is primarily linear. This is understood regarding existing results, which show that the elliptic flow $v_{2}$ is mainly proportional to the second cumulant of initial space anisotropy $\epsilon_{2}$ [69, 70, 71, 72, 68], except that the elliptic flow persists for ultra-central collisions [73, 74, 72, 75]. Nonetheless, the model could not faithfully reproduce the data at large transverse momenta. By tuning individual deformation parameters, the above results mostly hold. Specifically, the elliptic flow generally becomes more suppressed, particularly at the region with larger transverse momenta, as $\beta_{2}$ or $\beta_{4}$ increases. At smaller $p_{T}$, on the other hand, the dependence of elliptic flow on these parameters is complex. An explicit study on the dependence of $v_{2}$ on the deformation parameters will be presented below in Fig. 10, where the average flow will be analyzed.
A similar procedure is also carried out for various types of U+U collisions. In Fig. 8, we present the flow harmonics for the three Uranium nuclei while tuning the respective parameters regarding the quadrupole and hexadecapole deformations. The numerical results obtained by CHESS are also compared to the STAR data [66]. Generally, the initial geometric deformation plays a role in the final elliptic flow. As expected, the differential elliptic flow mainly increases with increasing $\beta_{2}$ or $\beta_{4}$, which is apparent, particularly for the region with large transverse momentum.
@@
Figure 8: Differential elliptic flow $v_{2}$ for U+U collisions at $\sqrt{s_{NN}}=193$ GeV with centrality 0-6% and different degrees of nucleus deformation. The numerical results obtained by the CHESS code are compared with the STAR data [66] for Au+Au collisions at 200 GeV. Top-Left: The results obtained by varying quadrupole deformation parameter $\beta_{2}$ for U. Top-Right: The results obtained by varying hexadecapole deformation parameter $\beta_{4}$ for U. Middle-Left: The results obtained by varying quadrupole deformation parameter for U2. Middle-Right: The results obtained by varying hexadecapole deformation parameter $\beta_{2}$ for U3. Bottom-left: The results obtained by varying hexadecapole deformation parameter $\beta_{4}$ for U3. Bottom-right: The results obtained for symmetric U+U, U2+U2, and U3+U3 collisions. The specific parameters are given in Tab. 1 for all the plots. The error bars indicate the standard errors.
@@
By comparing the results of Au+Au and U+U collisions, one seems to draw somewhat different observations about the effect of deformation on the collective flow. On the one hand, for U+U collisions, we have observed that the increase of deformation has, by and large, promoted the elliptic flow $v_{2}$, in accordance with Eq. (7). On the other hand, for Au+Au collisions, the observable effect is less substantial. In particular, at the high transverse momentum region, while the numerical results overshot the data, flow harmonics only increase slightly as the deformation becomes more significant. In other words, one finds that the elliptic flow in U+U collisions is more sensitive to the deformation. We attribute the difference in the magnitude of the effect to the highly nonlinear nature of the underlying hydrodynamic evolution and the strongly interacting system’s equation of state.
@@
Figure 9: Elliptic flow $v_{2}$ comparison for Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV and different U+U collisions at $\sqrt{s_{NN}}=193$ GeV. Top-Left: Symmetric Au nuclei compared to symmetric and deformed U nuclei in terms of $\beta_{2}$. Top-Right: Symmetric Au nuclei compared to symmetric and deformed U2 nuclei. Middle-Left: Symmetric Au nuclei compared to symmetric and deformed U3 nuclei in terms of $\beta_{2}$. For all the plots, the red triangles are used to represent the results for symmetric Au nuclei. Middle-Right: Symmetric Au nuclei compared to symmetric and deformed U nuclei in terms of $\beta_{4}$. Bottom: Symmetric Au nuclei compared to symmetric and deformed U3 nuclei in terms of $\beta_{4}$. The error bars indicate the standard errors.
@@
For Au+Au collisions, it is observed that the elliptic flow is rather sensitive when the deformation is small. When compared with Fig. 7, it seems that the somewhat outstanding value of $\bar{v_{2}}$ is primarily due to the contribution from the anisotropy of low $p_{T}$ particles. Here, the average is taken among all the emitted particles and on an event-by-event basis. As the deformation increases, the impact on the elliptic flow primarily follows a linear relation. The elliptic flow does not go to zero at vanishing $\beta_{2}$, which is largely attributed to non-vanishing geometric fluctuations. For the present study, we found that the value $\beta_{2}\approx-0.26$ gives the average elliptic flow that is in reasonable agreement with the experimental data. Additionally, regarding Fig. 10, the value $\beta_{4}\approx-0.031$ provides the best fit to the experimental data, which was originally adopted by $T_{R}ENT_{o}$. As a comparison, Giacalone [40] found a relatively smaller value $\beta_{2}\approx-0.15$, by considering only the quadrupole deformation. We attribute such a difference mostly to that of the hydrodynamic model and IC.
@@ IV Concluding remarks
Our numerical results prompt a comparison between Au+Au and U+U nuclear collisions, revealing subtle differences in the resulting behavior. In the case of Au+Au collisions, the simulations for both symmetric and deformed nuclei give reasonable results regarding the experimental data from the STAR Collaboration, as shown in Fig. 7. The results for deformed nuclei with slightly augmented flow harmonics stay closer to the experimental data. On the other hand, as shown in Fig. 10, the transition from symmetric to deformed nuclei marks a shift in average elliptic flow, $\overline{v_{2}}$, owing to the contributions from high transverse momentum $p_{T}\sim 2$ GeV. For symmetric nuclei, $\overline{v_{2}}$ significantly overestimates the data, and as deformation increases, the average elliptic flow converges to the experimental data. Conversely, for U+U nucleus collisions, numerical simulations diverge further from the STAR data as deformation factors increase. These results are explicitly illustrated in Fig. 8, where the elliptic flow increases in response to the deformation factors. The computation of the average elliptic flow, presented in Fig. 10, supports the linear behavior described in Eq. (7) that is extensively documented in the literature.
In the literature, the studies regarding the effect of initial geometric deformation are mainly motivated to answer the following question: whether such deformation is observable and which deformation factor, $\beta_{2}$ or $\beta_{4}$, predominantly influences collective flow? From the present study, the answer remains elusive, aligned with the results obtained by other authors [50, 38, 39, 40]. Nonetheless, for collisions of almost symmetric Au nuclei, the resulting integrated flow is found to be somewhat sensitive to the initial quadrupole and hexadecapole deformations. In this regard, it is interesting to scrutinize the topic by employing advanced analysis tools, such as Bayesian analysis [76, 77, 75] and maximum likelihood estimator [78, 79], and in particular emulator [80, 81, 82] and variational autoencoder, which have demonstrated their significant potential recently [83, 84, 85, 86].